If 0.2997 g of potassium iodate (KIO3) are dissolved in 25.00 mL of distilled water, what is the resulting molarity (M)? Give a brief explanation about why this is a very precise concentration to use. In other words, what makes potassium iodate such a good reagent with which to work in lab?
Answer #1

firstly, we calculate the molarity of potassium iodate.

since , Molarity = no. Of mole/ volume (in Litre)

So, number of moles of potassium iodate= weight in g / molar mass.

= 0.2997/214 = 0.0014

And volume = 25 ml/1000l = 0.025 Litre.

Therefore, Molarity M= 0.0014/0.025 = 0.056 mol/litre.

Now, Why makes potassium iodate such a good reagent ??

the answer is -

Since potassium iodate is a salt of iodine , and it is quite stable reagent.it has much higher shelf life. It can be stored in lab and can be used for a long time.

As well, potassium iodate is a thyroid blocking agent and also it can be used as a maturing agent in baking sometimes.

These unique properties makes potassium iodate a good reagent with which to work in lab.

please upvote